b+k
---= t solve for k
j
i know you have to get k by itself but i only got it down to bj+kj=jt, kj=-bj+jt so would it be k=-b+t??
i dont know if i did it right

Answer by Alan3354(30983)

I think you mean (b+k)/j = t
Multiply by j
b+k = jt
Subtract b
k = jt - b
------------